Hello im studying sets and logic and i need help with some problems im not getting.

1.Let U =(a,g,k,q,j) A=(a,q,j) B=(k,q) C=(a,j) find A-B
2.let n(A)=20, n(B)=35, and n(A∩B)=4. Find n(A∪B)

Answer by richard1234(7193)

A-B = {a, j} (elements in A that are not in B)
By inclusion-exclusion, |A∪B| = |A| + |B| - |A∩B| = 20 + 35 - 4 = 51.